OpenStudy (mathwizzard3):

A monument stands on level ground. The angle of elevation to the top of the monument taken at a point 405 feet away is 32 degrees. Find the height of the monument.

OpenStudy (mathwizzard3):

I need help with figuring out how to do this. ~Thanks

OpenStudy (isaidavila):

Tg32=h/405 Use that yo solve it

OpenStudy (jdoe0001):

 \(\bf sin(\theta)=\cfrac{opposite}{hypotenuse} \qquad \qquad % cosine cos(\theta)=\cfrac{adjacent}{hypotenuse} \\ \quad \\ % tangent \qquad \bullet\quad tan(\theta)=\cfrac{opposite}{adjacent} \\ \quad \\ \quad \\ tan(\theta)=\cfrac{opposite}{adjacent}\implies adjacent\cdot tan(\theta)=opposite\)